OpenStudy (anonymous):

Your next math test is worth 111 points and contains 32 problems. Each problem is worth either 4 points or 3 points. How many 3 point problems are on the test? A) 14 B) 15 Eliminate C) 17 D) 18

OpenStudy (anonymous):

@ajaylounsber

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

Okay, look at the choices. 14 15 17 18 If there were fourteen 3-point questions then that would be 3 \(\times\) 14 = 42 points.

OpenStudy (anonymous):

If you have 14 3-point questions. then you must have 18 4-point questions. 18 \(\times\) 4 = 72 72 + 42 = 114. So that's too much.

OpenStudy (anonymous):

So the answer is NOT 14

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let's try 15: 15 \(\times\) 3 = 45 32 - 15 = 17 17 \(\times\) 4 = 68 68 + 45 = 113. So the answer is NOT 15 either.

OpenStudy (anonymous):

18

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Let's try 17: 17 \(\times\) 3 = 51 32 - 17 = 15 15 \(\times\) 4 = 60 60 + 51 = 111 So your answer is 17. \(\huge\cal\rlap{\color{lime}-\color{lime}G\color{lime}r\color{lime}e\color{lime}e\color{lime }n\color{lime}ツ}{\color{black}{\;-Green~ツ}}\)

OpenStudy (anonymous):

thanks